# Rethinking Consistent Multi-Label Classification under Inexact Supervision

### v1

@@ Abstract
Partial multi-label learning and complementary multi-label learning are two popular weakly supervised multi-label classification paradigms that aim to alleviate the high annotation costs of collecting precisely annotated multi-label data. In partial multi-label learning, each instance is annotated with a candidate label set, among which only some labels are relevant; in complementary multi-label learning, each instance is annotated with complementary labels indicating the classes to which the instance does not belong. Existing consistent approaches for the two paradigms either require accurate estimation of the generation process of candidate or complementary labels or assume a uniform distribution to eliminate the estimation problem. However, both conditions are usually difficult to satisfy in real-world scenarios. In this paper, we propose consistent approaches that do not rely on the aforementioned conditions to handle both problems in a unified way. Specifically, we propose two unbiased risk estimators based on first- and second-order strategies. Theoretically, we prove consistency w.r.t. two widely used multi-label classification evaluation metrics and derive convergence rates for the estimation errors of the proposed risk estimators. Empirically, extensive experimental results validate the effectiveness of our proposed approaches against state-of-the-art methods.
@@ 1 Introduction
To address this, learning from weak supervision has become a prevailing way to mitigate the bottleneck of annotation cost for MLC (Sugiyama et al., 2022). Among them, partial multi-label learning (PML) and complementary multi-label learning (CML) have become two popular MLC paradigms. In PML, each instance is annotated with a candidate label set, among which only some labels are relevant but inaccessible to the learning algorithm (Xie and Huang, 2018; Sun et al., 2019; Gong et al., 2021). In CML, each instance is annotated with complementary labels, which indicate the classes to which the instance does not belong (Gao et al., 2023). Given that all relevant labels are included in the candidate label set, non-candidate labels contain no relevant labels and can be considered complementary labels, and vice versa. This suggests that the two problems are mathematically equivalent. Therefore, in this paper, we treat them as MLC under inexact supervision in a unified way. Figure 1 shows an example image annotated with inexact annotations. The label space contains ten labels in total. The candidate label set consists of four relevant labels {apple, plate, table, jug} and two false-positive ones {grapes, pear}. By excluding the candidate labels from the label space, the remaining four labels are {banana, cup, knife, flower}, which can be considered complementary labels. PML and CML do not require precise determination of all relevant labels during annotation, which demonstrates their great potential for alleviating annotation challenges in MLC.
@@
To this end, we propose a novel framework named COMES, i.e., COnsistent Multi-label classification under inExact Supervision. Based on a data generation process that does not use transition matrices, we introduce two instantiations with unbiased risk estimators w.r.t. the Hamming loss and ranking loss, respectively. Table 1 compares our approach with existing consistent PML and CML approaches. Our contributions are summarized as follows:
  • We propose a general and consistent framework for multi-label classification under inexact supervision that neither requires estimating the generation process of candidate or complementary labels nor relies on the uniform-distribution assumption.
  • We introduce risk-correction approaches to improve the generalization performance of the proposed unbiased risk estimators. We further prove consistency w.r.t. two widely used metrics and derive convergence rates of estimation errors for the proposed risk estimators.
@@ 2.1 Multi-Label Classification
where $\ell$ is a non-negative binary loss function, such as the binary cross-entropy loss. It is important to note that the Hamming loss only considers first-order model predictions and cannot account for label correlations. The ranking loss explicitly considers the ordering relationship between model outputs for a pair of labels. Specifically, the risk of $\bm{f}$ w.r.t. the ranking loss is [^2] To facilitate the analysis in this paper, we consider the coefficients of the losses for different label pairs to be 1 (Gao and Zhou, 2013; Xie and Huang, 2021, 2023).
@@ 3.1 Data Generation Process
The proof can be found in Appendix B.1. According to Lemma 1, the conditional density of instances where the $j$-th class is considered a non-candidate label is equivalent to the conditional density of instances where the $j$-th class is irrelevant. Notably, our data distribution assumption differs from both the uniform distribution assumption and the use of a transition matrix to flip the labels. Since the conditional probabilities of different candidate label sets can be different, our setting is more general than the uniform distribution assumption (Gao et al., 2023, 2025).
@@ 3.2 First-Order Strategy
Assume that there exists a constant $C_{\mathcal{G}}$ such that $\sup_{g_{j}\in\mathcal{G}}\|g\|_{\infty}\leq C_{\mathcal{G}}$ and a constant $C_{\ell}$ such that $\sup_{|z|\leq C_{\mathcal{G}}}\ell(z,y)\leq C_{\ell}$, where $\mathcal{G}$ is the model class. Since $\ell$ is non-negative, we assume that there exists a non-negative constant $\alpha$ such that $\forall j\in\mathcal{Y},\pi_{j}\mathbb{E}_{p(\bm{x}|y_{j}=1)}\left[\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]\geq\alpha$. Then, the bias of the expectation of the corrected risk estimator w.r.t. the $\ell$-risk w.r.t. the Hamming loss has the following lower and upper bounds:
@@ 3.3 Second-Order Strategy
Since $R_{\mathrm{R}}^{\ell}(\bm{g})>0$, we assume that there exists a non-negative constant $\gamma$ such that $R_{\mathrm{R}}^{\ell}(\bm{g})\geq\gamma$. We also assume that $\beta$ is chosen such that $\beta\leq\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)$. By the assumptions in Theorem 2, the bias of the expectation of the corrected risk estimator w.r.t. the ranking loss has the following lower and upper bounds:
@@ 4.3 Sensitivity Analysis
Influence of $\beta$. We also investigated the influence of the hyperparameter $\beta$ for the flooding regularization used in COMES-RL. From Figure 3, we observe that the performance of COMES-RL is rather stable when $\beta$ is set within a reasonable range. During our experiments, we found that the performance was already competitive by setting $\beta=0$ on many datasets. However, the performance may degrade when $\beta$ is set to a large value. This also matches our theoretical results in Theorem 5, where consistency holds when $\beta$ is not too large.
@@ 5 Conclusion
In this paper, we rethought MLC under inexact supervision by proposing a novel framework. We proposed two instantiations of unbiased risk estimators w.r.t. the Hamming loss and ranking loss, two widely used evaluation metrics for MLC, respectively. We also introduced risk-correction approaches to improve generalization performance with theoretical guarantees. Extensive experiments on ten real-world and synthetic benchmark datasets validated the effectiveness of the proposed approaches. A limitation of this work is that we consider the generation process to be independent of the instances. In the future, it is promising to extend our proposed methodologies to instance-dependent settings.
@@ A.1 Algorithmic Pseudo-Code
8:   else if using the COMES-HL algorithm then
@@ A.2 Class Prior Estimation
We can use any off-the-shelf mixture proportion estimation algorithm to estimate the class priors [Ramaswamy et al., 2016, Garg et al., 2021, Yao et al., 2022, Zhu et al., 2023], which are mainly designed to estimate the class prior with positive and unlabeled data for binary classification. Specifically, we generate negative and unlabeled datasets according to Eq. (6) and then apply the mixture proportion estimation algorithm. The algorithmic details are summarized in Algorithm 2.
@@ B.1 Proof of Lemma 1
On one hand, we have
@@ B.4 Proof of Theorem 3
Lemma 4 (Theorem 4.12 in [Ledoux and Talagrand, 1991]).
@@
where the inequalities are due to the triangle inequality. Then, by combining Inequalities (22) and (23), it is a routine work [Mohri et al., 2012] to show that
@@ B.5 Proof of Corollary 1
Lemma 6 (Theorem 4 in Gao and Zhou [2013]).
@@
Lemma 7 (Theorem 32 in Gao and Zhou [2013]).
@@ B.6 Proof of Theorem 4
$\left.-\int\left(\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)+\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right)p\left(\bm{x},y_{j}=0,y_{k}=0\right)\mathop{}\!\mathrm{d}\bm{x}\right)$
@@
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|y_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Cp(y_{j}=0,y_{k}=0)\right).$
@@
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|s_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Cp(y_{j}=0,y_{k}=0)\right),$
@@ B.8 Proof of Theorem 6
where the first inequality is by symmetrization [Mohri et al., 2012] and the second inequality is by Lemma 4. The proof is complete. ∎
@@
where the inequalities are due to the triangle inequality. Then, by combining Inequalities 30 and 31, it is a routine work [Mohri et al., 2012] to show that
@@ B.9 Proof of Corollary 2
Lemma 11 (Theorem 10 in Gao and Zhou [2013]).
@@ Appendix C Details of Experiments
C.1 Synthetic Dataset Generation
For synthetic datasets, we generated partial labels in two ways. In Case-a, we flipped negative labels to positive labels independently with probability $\tau$. In Case-b, for each class we deterministically enforced a flip rate of $\rho$ by uniformly sampling a $\rho$ fraction of negative instances in that class and flipping them to positive. We set $\tau{=}\rho{=}0.9$ in this paper.
C.2 Baseline
We evaluate against five classical baselines commonly used in PML/CML learning. (A) BCE: uses the given partial label as the cross-entropy target. (B) CCMN [Xie and Huang, 2023]: treats PML as multi-label classification with class-conditional noise, relying on a noise transition matrix. (C) GDF [Gao et al., 2023]: proposes an unbiased risk estimator for multi-labeled single complementary label learning. (D) CTL [Gao et al., 2025]: introduces a risk-consistent approach by rewriting the loss function. (E) MLCL [Gao et al., 2024]: estimates an initial transition matrix via binary decompositions, then refines it with label correlations.
C.3 Implementation Details
For real-world datasets, we used an MLP encoder for all baselines, trained for 200 epochs with a learning rate of 5e-3, weight decay of 1e-4, and the SGD optimizer with cosine decay. For synthetic image datasets, we adopted a ResNet-50 backbone pretrained on ImageNet [Deng et al., 2009], trained for 30 epochs with a learning rate of 1e-4 using the Adam optimizer. We used the binary cross-entropy loss. For fair comparisons, we used the same setup across all baselines. We assumed that the class priors were accessible to the learning algorithm.

### v2

@@ Abstract
Partial multi-label learning and complementary multi-label learning are two popular weakly supervised multi-label classification paradigms that aim to alleviate the high annotation costs of collecting precisely annotated multi-label data. In partial multi-label learning, each instance is annotated with a candidate label set, among which only some labels are relevant; in complementary multi-label learning, each instance is annotated with complementary labels indicating the classes to which the instance does not belong. Existing consistent approaches for the two paradigms either require accurate estimation of the generation process of candidate or complementary labels or assume a uniform distribution to eliminate the estimation problem. However, both conditions are usually difficult to satisfy in real-world scenarios. In this paper, we propose consistent approaches that do not rely on the aforementioned conditions to handle both problems in a unified way. Specifically, we propose two risk estimators based on first- and second-order strategies. Theoretically, we prove consistency w.r.t. two widely used multi-label classification evaluation metrics and derive convergence rates for the estimation errors of the proposed risk estimators. Empirically, extensive experimental results on both real-world and synthetic datasets validate the effectiveness of our proposed approaches against state-of-the-art methods.
@@ 1 Introduction
To address this, learning from weak supervision has become a prevailing way to mitigate the bottleneck of annotation cost for MLC (Sugiyama et al., 2022). Among them, partial multi-label learning (PML) and complementary multi-label learning (CML) have become two popular MLC paradigms. In PML, each instance is annotated with a candidate label set, among which only some labels are relevant but inaccessible to the learning algorithm (Xie and Huang, 2018; Sun et al., 2019; Gong et al., 2021). In CML, each instance is annotated with complementary labels, which indicate the classes to which the instance does not belong (Gao et al., 2023; Zhu et al., 2025; Gao et al., 2026). Given that all relevant labels are included in the candidate label set, non-candidate labels contain no relevant labels and can be considered complementary labels, and vice versa. This suggests that the two problems are mathematically equivalent. Therefore, in this paper, we treat them as MLC under inexact supervision in a unified way. Figure 1 shows an example image annotated with inexact annotations. The label space contains ten labels in total. The candidate label set consists of four relevant labels {apple, plate, table, jug} and two false-positive ones {grapes, pear}. By excluding the candidate labels from the label space, the remaining four labels are {banana, cup, knife, flower}, which can be considered complementary labels. PML and CML do not require precise determination of all relevant labels during annotation, which demonstrates their great potential for alleviating annotation challenges in MLC.
@@
To this end, we propose a novel framework named COMES, i.e., COnsistent Multi-label classification under inExact Supervision. Based on a data generation process that does not use transition matrices, we introduce two instantiations with risk estimators w.r.t. the Hamming loss and ranking loss, respectively. Table 1 compares our approach with existing consistent PML and CML approaches. Our contributions are summarized as follows:
  • We propose a consistent framework for multi-label classification under inexact supervision that neither requires estimating the generation process of candidate or complementary labels nor relies on the uniform-distribution assumption.
  • We introduce risk-correction approaches to improve the generalization performance of the proposed risk estimators. We further prove consistency w.r.t. two widely used metrics and derive convergence rates of estimation errors for the proposed risk estimators.
@@ 2.1 Multi-Label Classification
where $\ell$ is a non-negative binary loss function, such as the binary cross-entropy loss. It is important to note that the Hamming loss only considers first-order model predictions and cannot account for label correlations. The ranking loss explicitly considers the ordering relationship between model outputs for a pair of labels. Specifically, the risk of $\bm{f}$ w.r.t. the ranking loss is [^2] To facilitate the analysis in this paper, we consider the coefficients of the losses for different label pairs to be 1 (Gao and Zhou, 2013; Xie and Huang, 2021; 2023).
@@ 3.1 Data Generation Process
The proof can be found in Appendix B.1. According to Lemma 1, the conditional density of instances where the $j$-th class is considered a non-candidate label is equivalent to the conditional density of instances where the $j$-th class is irrelevant. Notably, our data distribution assumption differs from both the uniform distribution assumption and the use of a transition matrix to flip the labels. Since the conditional probabilities of different candidate label sets can be different, our setting is more general than the uniform distribution assumption (Gao et al., 2023; 2025).
@@ 3.2 First-Order Strategy
Assume that there exists a constant $C_{\mathcal{G}}$ such that $\sup_{g_{j}\in\mathcal{G}}\|{g_{j}}\|_{\infty}\leq C_{\mathcal{G}}$ and a constant $C_{\ell}$ such that $\sup_{|z|\leq C_{\mathcal{G}}}\ell(z,y)\leq C_{\ell}$, where $\mathcal{G}$ is the model class. We assume that there exists a positive constant $\alpha$ such that $\forall j\in\mathcal{Y},\pi_{j}\mathbb{E}_{p(\bm{x}|y_{j}=1)}\left[\ell\left(g_{j}\left(\bm{x}\right),1\right)\right]\geq\alpha$. Then, the bias of the expectation of the corrected risk estimator w.r.t. the $\ell$-risk w.r.t. the Hamming loss has the following lower and upper bounds:
@@ 3.3 Second-Order Strategy
We assume that there exists a positive constant $\gamma$ such that $R_{\mathrm{R}}^{\ell}(\bm{g})\geq\gamma$. We also assume that $\beta$ is chosen such that $\beta\leq\sum\nolimits_{1\leq j<k\leq q}Mp(y_{j}=0,y_{k}=0)z$. By the assumptions in Theorem 2, the bias of the expectation of the corrected risk estimator w.r.t. the ranking loss has the following lower and upper bounds:
@@ 4.1 Experimental Setup
For real-world datasets, we used an MLP encoder for all baselines, trained for 200 epochs with a learning rate of 5e-3, weight decay of 1e-4, and the SGD optimizer with cosine decay. For synthetic image datasets, we adopted a ResNet-50 backbone pretrained on ImageNet (Deng et al., 2009), trained for 30 epochs with a learning rate of 1e-4 using the Adam optimizer. For fair comparisons, we used the same setup across all baselines. We assumed that the class priors were accessible to the learning algorithm. We instantiated $\ell$ with the binary cross-entropy loss for COMES-HL and the sigmoid loss for COMES-RL. We performed ten-fold cross-validation on real-world datasets. This means we used nine folds for training and one fold for testing. Then, we recorded the mean accuracy and standard deviation. For the synthetic datasets, we generated synthetic labels three times and recorded the mean accuracy and standard deviation. Finally, we conducted paired t-tests at a 0.05 significance level.
@@ 4.3 Sensitivity Analysis
Influence of $\beta$. We also investigated the influence of the hyperparameter $\beta$ for the flooding regularization used in COMES-RL. From Figure 3, we observe that the performance of COMES-RL is rather stable when $\beta$ is set within a reasonable range on the mirflickr dataset. During our experiments, we found that the performance was already competitive by setting $\beta=0$ on many datasets. However, the performance may degrade when $\beta$ is set to a large value. This also matches our theoretical results in Theorem 5, where consistency holds when $\beta$ is not too large.
@@ 5 Conclusion
In this paper, we rethought MLC under inexact supervision by proposing a novel framework. We proposed two instantiations of risk estimators w.r.t. the Hamming loss and ranking loss, two widely used evaluation metrics for MLC, respectively. We also introduced risk-correction approaches to improve generalization performance with theoretical guarantees. Extensive experiments on ten real-world and synthetic benchmark datasets validated the effectiveness of the proposed approaches. A limitation of this work is that we consider the generation process to be independent of the instances. In the future, it is promising to extend our proposed methodologies to instance-dependent settings.
Acknowledgments
WW was supported by the Junior Research Associate (JRA) program of RIKEN. MS was supported by JST ASPIRE Grant Number JPMJAP2405.
@@ A.1 Algorithmic Pseudo-Code
8:   else if using the COMES-RL algorithm then
@@ A.2 Class Prior Estimation
We can use any off-the-shelf mixture proportion estimation algorithm to estimate the class priors (Ramaswamy et al., 2016; Garg et al., 2021; Yao et al., 2022; Zhu et al., 2023), which are mainly designed to estimate the class prior with positive and unlabeled data for binary classification. Specifically, we generate negative and unlabeled datasets according to Eq. (6) and then apply the mixture proportion estimation algorithm. The algorithmic details are summarized in Algorithm 2.
@@ B.1 Proof of Lemma 1
According to the definition of PML, when $s_{j}=0$, the $j$-th class is impossible to be a relavant label, and we have $p(j\notin Y|\bm{x},s_{j}=0)=p(j\notin Y|s_{j}=0)=1$. Therefore, on one hand, we have
@@ B.4 Proof of Theorem 3
Lemma 4 (Theorem 4.12 in (Ledoux and Talagrand, 1991)).
@@
where the inequalities are due to the triangle inequality. Then, by combining Inequalities (22) and (23), it is a routine work (Mohri et al., 2012) to show that
@@ B.5 Proof of Corollary 1
Lemma 6 (Theorem 4 in Gao and Zhou (2013)).
@@
Lemma 7 (Theorem 32 in Gao and Zhou (2013)).
@@ B.6 Proof of Theorem 4
$\left.-\int\left(\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),0\right)+\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right)p\left(\bm{x},y_{j}=0,y_{k}=0\right)\mathop{}\!\mathrm{d}\bm{x}\right)$
@@
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|y_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Mp(y_{j}=0,y_{k}=0)\right).$
@@
$\left.+(1-\pi_{k})\mathbb{E}_{p(\bm{x}|s_{k}=0)}\left[\ell\left(g_{j}(\bm{x})-g_{k}(\bm{x}),1\right)\right]-Mp(y_{j}=0,y_{k}=0)\right),$
@@ B.8 Proof of Theorem 6
where the first inequality is by symmetrization (Mohri et al., 2012) and the second inequality is by Lemma 4. The proof is complete. ∎
@@
where the inequalities are due to the triangle inequality. Then, by combining Inequalities 30 and 31, it is a routine work (Mohri et al., 2012) to show that
@@ B.9 Proof of Corollary 2
Lemma 11 (Theorem 10 in Gao and Zhou (2013)).
@@ Appendix C Details of Experiments
C.1 More Details of Datasets
For synthetic datasets, we consider two data generation processes. In case-a, irrelevant labels are flipped to candidate labels independently, which is the assumption used in Xie and Huang (2023). This strategy is common in learning with noisy labels (Han et al., 2018), where PML is a special case of MLC with noisy labels (Xie and Huang, 2023). In case-b, we assign non-candidate labels in a class-wise manner. For each class, we randomly sample a fraction of the training data and assign that class as a non-candidate label. This data generation process corresponds to the assumption proposed in this paper. We use this process to confirm the effectiveness of our proposed method under this assumption. Additionally, we selected high flipping rates to evaluate the effectiveness of our proposed methods on challenging datasets with high noise rates since real-world datasets have low noise rates. In this paper, we set the flipping rate in Case-a to be 0.9 and and the sampling rate in Case-b to be 0.1.
C.2 Baselines
We evaluate against five classical baselines commonly used in PML/CML learning. (A) BCE: uses the given candidate label as the cross-entropy target. (B) CCMN (Xie and Huang, 2023): treats PML as multi-label classification with class-conditional noise, relying on a noise transition matrix. (C) GDF (Gao et al., 2023): proposes an unbiased risk estimator for multi-labeled single complementary label learning. (D) CTL (Gao et al., 2025): introduces a risk-consistent approach by rewriting the loss function. (E) MLCL (Gao et al., 2024): estimates an initial transition matrix via binary decompositions, then refines it with label correlations.
Appendix D Instance-Dependent Cases
The current literature on partial multi-label learning (PML) and complementary multi-label learning (CML) assumes that label generation is independent of instances (see Table 1). Following previous work, we also consider the instance-independent case. It is very challenging to design consistent methods for instance-dependent cases due to the difficulty of estimating instance-dependent generation processes, as far as we know from the literature on weakly supervised learning. In future work, we will consider developing instance-dependent methods with strong theoretical guarantees.